# On the magnetic perturbation theory for Chern insulators

### v1

@@ 5. Lack of continuity in the norm topology
Finally, it is worth mentioning that magnetic spectral projections might fail to be norm continuous even in the topologically trivial case, i.e. when their Chern character is equal to zero, as it is shown in [17, Corollary 1.2] by using the concept of exponentially localized magnetic Wannier functions.
@@ Appendix A Continuity and off-diagonal decay
In this section we collect three more general technical lemmas that we used in the previous proofs. Lemma A.1 and A.2 deal with the continuity and exponential localization of the integral kernel of an operator power series. They are a generalization of the result contained in [17, Appendix C]. Moreover, Lemma A.3 provides an estimate on the difference of local traces for projections that can be intertwined by a Kato-Nagy unitary.
We start with a generalization of the result of Lemma 3.5. Instead of requiring a uniform control on $\epsilon$ dependence of the integral kernel of the operator $D_{\epsilon}$, we only require the exponential localization of the integral kernel and the control on the smallness of the operator norm.
@@
$|\big{(}F(D)\big{)}({\bf x};{\bf y})|\leq C^{\prime}\,e^{-\alpha^{\prime}\|{\bf x}-{\bf y}\|}\,\quad\forall\,{\bf x}\,,{\bf y}\in{\mathbb{R}}^{2}.$
@@
We first show that there exists $C>0$ such that for all ${\bf x}_{0}\in{\mathbb{R}}^{2}$ and $20<\alpha^{\prime}<\alpha/$ 2 we have
$\|e^{\alpha^{\prime}\|\cdot-{\bf x}_{0}\|}De^{-\alpha^{\prime}\|\cdot-{\bf x}_{0}\|}-D\|\leq C\alpha^{\prime}.$ (A.1)
@@
and the Schur test finishes the proof of (A.1). Thus if $\alpha^{\prime}$ is small enough then
@@
are bounded operators. Moreover, $DG(D)D$ is an integral operator with a kernel given by
@@
We apply Lemma 3.1 to $D\chi_{\Omega_{n}}G(D)\chi_{\Omega_{n}}D$ and get that the scalar products $\langle\overline{D({\bf x};\cdot)},\chi_{\Omega_{n}}G(D)\chi_{\Omega_{n}}D(\cdot;{\bf y})\rangle$ are jointly continuous functions, hence $\langle\overline{D({\bf x};\cdot)},G(D)D(\cdot;{\bf y})\rangle$ is also jointly continuous as a uniform limit.
@@
$\langle\overline{D({\bf x};\cdot)},G(D)D(\cdot;{\bf y})\rangle=\langle\overline{D({\bf x};\cdot)}e^{-\alpha^{\prime}\|\cdot-{\bf y}\|},G\big{(}e^{\alpha^{\prime}\|\cdot-{\bf y}\|}De^{-\alpha^{\prime}\|\cdot-{\bf y}\|}\big{)}e^{\alpha^{\prime}\|\cdot-{\bf y}\|}D(\cdot,{\bf y})\rangle$
@@
$U$ $=\left[{\bf 1}-\left(P_{1}-P_{2}\right)^{2}\right]^{-1/2}\big{(}P_{1}P_{2}+\left({\bf 1}-P_{1}\right)\left({\bf 1}-P_{2}\right)\big{)}$ (A.2)
and for which $P_{1}U=UP_{2}$, has the property that $U-{\bf 1}$ has a jointly continuous integral kernel and there exists $C^{\prime}>0$ and $0<\alpha^{\prime}<\alpha$ such that
@@
[17, Lemma 2.1] Let $P_{1}$, $P_{2}$ and $U$ as in Lemma A.2. For $L\geq 1$, let $\chi_{L}$ be the characteristic function of the set $\Lambda_{L}:=[-L,L]^{2}$. Then $\chi_{L}P_{j}$ are trace class, and there exists another constant $C_{2}<\infty$ such that
@@
Exploiting the invariance under unitary conjugation of the trace we get the equality $\operatorname{Tr}(\chi_{L}P_{2}U^{*}U)=\operatorname{Tr}(U\chi_{L}P_{2}U^{*})$, thus
@@
$\int_{{\bf x}\in{\mathbb{R}}^{2}}\int_{{{\bf x}^{\prime}}\in{\mathbb{R}}^{2}}e^{-{\alpha^{\prime}}\|{\bf x}-{\bf x}^{\prime}\|}|\chi_{L}({\bf x})-\chi_{L}({\bf x}^{\prime})|\mathrm{d}{\bf x}^{\prime}\mathrm{d}{\bf x}\;.$
In the integral (A), the integrand is non-zero only if one variable belongs to $\Lambda_{L}$ and the other one lies outside $\Lambda_{L}$. Then, due to symmetry, it is enough to estimate the integral
@@ Appendix B Integral kernel estimates for the resolvent
In this section we briefly show how to prove the estimates (1.5) and (1.6). Let us recall our assumptions on the scalar and magnetic potentials: we assume that the scalar potential $V$ is uniformly in $L^{2+\delta}_{\rm loc}({\mathbb{R}}^{2})$, for $\delta>0$, namely
$\|V\|_{2+\delta,\textrm{unif}}:=\sup_{{\bf y}\in{\mathbb{R}}^{2}}\int_{\|{\bf x}-{\bf y}\|\leq 1}|V({\bf x})|^{2+\delta}\mathrm{d}{\bf x}<+\infty\,,$ (B.1)
@@
With this assumption, it is a standard result that $V$ is infinitesimally operator bounded with respect to the free Laplacian [40]. Moreover, since the magnetic potential is smooth, we have that $H_{b}$ satisfies the diamagnetic inequality [1], this means that $V$ is also infinitesimally operator bounded with respect to the magnetic Laplacian $(-\mathrm{i}\nabla-bA-\mathcal{A})^{2}$. From [8, 43] (see also [37, Proposition 5.5]) we get that the resolvent of $H_{b}$ has an integral kernel that is jointly continuous outside of the diagonal. Without entering into the details of the argument, the main idea is to analyze the Schrödinger semigroup, namely the operator $e^{-tH_{b}}$, for $t>0$, and then, for $\lambda>0$ large enough, to use the Laplace transform:
@@
The key observation lies in the fact that the diamagnetic inquality allows a pointwise upper bound of the Schrödinger semigroup of $H_{b}$ provided by the Schrödinger semigroup of the free Laplacian, which is explicitly known. This leads to (1.5) with $z$ replaced by $-\lambda$. The estimate for an arbitrary $z$ is then obtained by repeatedly using the first resolvent identity.
It remains to show (1.6), which is a bit more delicate. First, notice that for smooth and bounded magnetic fields, the same estimate has been proved in [10, Proposition B.9], in particular for the purely magnetic operator $h_{b}:=(-\mathrm{i}\nabla-bA-\mathcal{A})^{2}$. From the second resolvent identity we get
@@
satisfies (1.6). In view of (1.5), it is enough to show the estimate (1.6) is satisfied by the following integral
@@
In order to exploit the fact that $V$ is uniformly locally in $L^{2+\delta}({\mathbb{R}}^{2})$, we start by bounding the integral $\mathfrak{I}$ by
@@
$|\mathfrak{I}_{\gamma}({\bf x},{\bf x}^{\prime})|$ $\leq C\,e^{-\alpha\|{\bf x}-{\bf x}^{\prime}\|/4}e^{-\alpha\|{\bf x}-\gamma\|/4}$ (B.2)
×$\qquad\times\int_{\|{\bf y}-\gamma\|\leq 1}\|{\bf x}-{\bf y}\|^{-1}|V({\bf y})|\,\ln(2+\|{\bf y}-{\bf x}^{\prime}\|^{-1})\,\mathrm{d}{\bf y}.$
@@ Acknowledgments
H.D.C. gratefully acknowledges the financial support from Grant 2032-00005B of the Danish Council for Independent Research |||| Natural Sciences. M.M. gratefully acknowledges the support of PNRR Italia Domani and Next Generation EU through the ICSC National Research Centre for High Performance Computing, Big Data and Quantum Computing and the support of the MUR grant Dipartimento di Eccellenza 2023–2027.

### v2

@@ 5. Lack of continuity in the norm topology
Finally, it is worth mentioning that magnetic spectral projections might fail to be norm continuous even in the topologically trivial case, i.e. when their Chern character is equal to zero, as it is shown in [15, Corollary 1.2] (see also the discussions in [37, 13]) by using the concept of exponentially localized magnetic Wannier functions.
@@ Appendix A Continuity and off-diagonal decay
In this section we collect three more general technical lemmas that we used in the previous proofs. Lemma A.1 and A.2 deal with the continuity and exponential localization of the integral kernel of an operator power series. They are a generalization of the result contained in [15, Appendix C]. Moreover, Lemma A.3 provides an estimate on the difference of local traces for projections that can be intertwined by a Kato-Nagy unitary.
We start with a generalization of the result of Lemma 3.5. Instead of requiring a uniform control on the $\epsilon$ dependence of the integral kernel of the operator $D_{\epsilon}$, we only require the exponential localization of the integral kernel and the control on the smallness of the operator norm.
@@
$\left|\big(F(D)\big)({\bf x};{\bf y})\right|\leq C^{\prime}\,e^{-\alpha^{\prime}\|{\bf x}-{\bf y}\|}\,\quad\forall\,{\bf x}\,,{\bf y}\in{\mathbb{R}}^{2}.$
@@
We first show that there exists $C>0$ such that for all ${\bf x}_{0}\in{\mathbb{R}}^{2}$ and $0<\alpha^{\prime}<\alpha/2$ we have
(A.1) $\|e^{\alpha^{\prime}\|\cdot-{\bf x}_{0}\|}De^{-\alpha^{\prime}\|\cdot-{\bf x}_{0}\|}-D\|\leq C\alpha^{\prime}.$
@@
and the Schur test [27, Theorem 5.2] finishes the proof of (A.1). Thus, if $\alpha^{\prime}$ is small enough then
@@
are bounded operators. Moreover, $DG(D)D$ is an integral operator with an integral kernel given by
@@
We apply Lemma 3.1 to $D\chi_{\Omega_{n}}G(D)\chi_{\Omega_{n}}D$ and we get that the scalar products
$\langle\overline{D({\bf x};\cdot)},\chi_{\Omega_{n}}G(D)\chi_{\Omega_{n}}D(\cdot;{\bf y})\rangle$
define jointly continuous functions, hence $\langle\overline{D({\bf x};\cdot)},G(D)D(\cdot;{\bf y})\rangle$ is also jointly continuous as a uniform limit.
@@
$\langle\overline{D({\bf x};\cdot)},G(D)D(\cdot;{\bf y})\rangle=\langle\overline{D({\bf x};\cdot)}e^{-\alpha^{\prime}\|\cdot-{\bf y}\|},G\big(e^{\alpha^{\prime}\|\cdot-{\bf y}\|}De^{-\alpha^{\prime}\|\cdot-{\bf y}\|}\big)e^{\alpha^{\prime}\|\cdot-{\bf y}\|}D(\cdot;{\bf y})\rangle$
@@
(A.2) $U$ $=\left[{\bf 1}-\left(P_{1}-P_{2}\right)^{2}\right]^{-1/2}\big(P_{1}P_{2}+\left({\bf 1}-P_{1}\right)\left({\bf 1}-P_{2}\right)\big)$
and for which $P_{1}U=UP_{2}$, has the property that $U-{\bf 1}$ has a jointly continuous integral kernel and there exist $C^{\prime}>0$ and $0<\alpha^{\prime}<\alpha$ such that
@@
[15, Lemma 2.1] Let $P_{1}$, $P_{2}$ and $U$ as in Lemma A.2. For $L\geq 1$, let $\chi_{L}$ be the characteristic function of the set $\Lambda_{L}:=[-L,L]^{2}$. Then $\chi_{L}P_{j}$, $j\in\{1,2\}$, are trace class, and there exists another constant $C_{2}<\infty$ such that
@@
By exploiting the invariance under unitary conjugation of the trace, we obtain the equality $\operatorname{Tr}(\chi_{L}P_{2}U^{*}U)=\operatorname{Tr}(U\chi_{L}P_{2}U^{*})$, thus
@@
(A.3) $\int_{{\bf x}\in{\mathbb{R}}^{2}}\int_{{{\bf x}^{\prime}}\in{\mathbb{R}}^{2}}e^{-{\alpha^{\prime}}\|{\bf x}-{\bf x}^{\prime}\|}|\chi_{L}({\bf x})-\chi_{L}({\bf x}^{\prime})|\mathrm{d}{\bf x}^{\prime}\mathrm{d}{\bf x}\;.$
In the integral (A.3), the integrand is non-zero only if one variable belongs to $\Lambda_{L}$ and the other one lies outside $\Lambda_{L}$. Then, due to symmetry, it is enough to estimate the integral
@@ Appendix B Integral kernel estimates for the resolvent
In this section we briefly show how to prove the estimates (1.4) and (1.5). Let us recall our assumptions on the scalar and magnetic potentials: we assume that the scalar potential $V$ is in $L^{2+\delta}_{\mathrm{u.loc}}({\mathbb{R}}^{2})$, for $\delta>0$, which means that it is uniformly in $L^{2+\delta}_{\rm loc}({\mathbb{R}}^{2})$, namely
(B.1) $\|V\|_{2+\delta,\textrm{unif}}:=\sup_{{\bf y}\in{\mathbb{R}}^{2}}\int_{\|{\bf x}-{\bf y}\|\leq 1}|V({\bf x})|^{2+\delta}\mathrm{d}{\bf x}<+\infty\,$,
@@
With this assumption, it is a standard result that $V$ is infinitesimally operator bounded with respect to the free Laplacian [41]. Moreover, since the magnetic potential is smooth, we have that $H_{b}$ satisfies the diamagnetic inequality [2], this means that $V$ is also infinitesimally operator bounded with respect to the magnetic Laplacian $(-\mathrm{i}\nabla-bA-\mathcal{A})^{2}$. From [9, 44] (see also [39, Proposition 5.5]) we get that the resolvent of $H_{b}$ has an integral kernel that is jointly continuous outside of the diagonal. Without entering into the details of the argument, the main idea is to analyze the Schrödinger semigroup, namely the operator $e^{-tH_{b}}$, for $t>0$, and then, for $\lambda>0$ large enough, to use the Laplace transform:
@@
The key observation lies in the fact that the diamagnetic inquality allows a pointwise upper bound of the Schrödinger semigroup of $H_{b}$ provided by the Schrödinger semigroup of the free Laplacian, which is explicitly known. This leads to (1.4) with $z$ replaced by $\lambda$. The estimate for an arbitrary $z$ is then obtained by repeatedly using the first resolvent identity.
It remains to show (1.5), which is a bit more delicate. First, notice that for smooth and bounded magnetic fields, the same estimate has been proved in [11, Proposition B.9], in particular for the purely magnetic operator $h_{b}:=(-\mathrm{i}\nabla-bA-\mathcal{A})^{2}$. From the second resolvent identity we get
@@
satisfies (1.5). In view of (1.4), it is enough to show that the estimate (1.5) is satisfied by the following integral
@@
In order to exploit the fact that $V$ is uniformly locally in $L^{2+\delta}({\mathbb{R}}^{2})$, we start by bounding the integral $\mathfrak{I}({\bf x},{\bf x}^{\prime})$ by
@@
(B.2) $\left|\mathfrak{I}_{\gamma}({\bf x},{\bf x}^{\prime})\right|$ $\leq C\,e^{-\alpha\|{\bf x}-{\bf x}^{\prime}\|/4}e^{-\alpha\|{\bf x}-\gamma\|/4}$
$\qquad\times\int_{\|{\bf y}-\gamma\|\leq 1}\|{\bf x}-{\bf y}\|^{-1}|V({\bf y})|\,\ln(2+\|{\bf y}-{\bf x}^{\prime}\|^{-1})\,\mathrm{d}{\bf y}.$
@@ Acknowledgments
H.D.C. gratefully acknowledges the financial support from Grant 2032-00005B of the Danish Council for Independent Research $|$ Natural Sciences. M.M. gratefully acknowledges the support of PNRR Italia Domani and Next Generation EU through the ICSC National Research Centre for High Performance Computing, Big Data and Quantum Computing and the support of the MUR grant Dipartimento di Eccellenza 2023–2027.